### v1

# Hierarchical Large Scale Multirobot Path (Re)Planning

Abstract
We consider a large-scale multi-robot path planning problem in a cluttered environment. Our approach achieves real-time replanning by dividing the workspace into cells and utilizing a hierarchical planner. Specifically, multi-commodity flow-based high-level planners route robots through the cells to reduce congestion, while an anytime low-level planner computes collision-free paths for robots within each cell in parallel. Despite resulting in longer paths compared to the baseline multi-agent pathfinding algorithm, our method produces a solution with significant improvement in computation time. Specifically, we show empirical results of a $500$-times speedup in computation time compared to the baseline multi-agent pathfinding approach on the environments we study. We account for the robot’s embodiment and support non-stop execution when replanning continuously. We demonstrate the real-time performance of our algorithm with up to $142$ robots in simulation, and a representative $32$ physical Crazyflie nano-quadrotor experiment.
I Introduction
Large fleets of robots, such as those used in warehouse operations [1], disaster response [2], and delivery [3], demand coordination solutions that adjust in real time to changing goals. In this work, we present a real-time lifelong hierarchical method for navigating a large team of robots to independent goals in a large, cluttered environment that guarantees collision avoidance. By lifelong, we mean robots can enter and exit the space, and can receive another goal at any time, as they would in a warehouse or delivery problem. Our approach partitions the space into disjoint cells, allowing planning algorithms to run concurrently in parallel within each cell. A high-level planner routes robots through the partition, while a low-level anytime multi-agent pathfinding (MAPF) algorithm navigates robots to local goals within each cell in parallel. The real-time property holds as long as there are not too many cells or robots in a workspace; the limits for real-time operation are empirical and problem-specific, however, we demonstrate real-time performance for 142 robots in simulation with a 25-cell partition.
Figure 1: Long exposure of $32$ quadrotors navigating a cluttered environment.
We are particularly interested in unmanned aerial vehicles (UAVs) operating in 3D space, such as in city-scale on-demand UAV package delivery, however, our approach applies to robots operating in 2D as well. We present two approaches for high-level planning depending on the problem’s requirements: 1) an egocentric greedy approach that always operates in real-time and 2) a novel high-level planner that routes robots through the partition using multi-commodity flow (MCF) [4]. There are tradeoffs between these two approaches. The egocentric greedy planner operates in real-time regardless of the number of cells; however, it has no mechanism for distributing robots, thus it can result in congestion and longer low-level planning times within some cells. On the other hand, the MCF-based approach eases cell congestion by regulating the flow of robots into each cell while ensuring bounded-suboptimal inter-cell routing; thus, it can be useful in environments such as urban UAV package delivery, where different types of cells (e.g., residential vs. highway) may have different limits on the influx of robots. The MCF-based planners can operate in real-time under certain conditions, thus allowing for lifelong replanning while reducing congestion, which leads to faster, real-time low-level planning within each cell.
The low-level planner ensures collision avoidance while respecting the robots’ geometric shape. A cell-crossing protocol allows robots to transition between cells without stopping in midair. Combined with the MCF-based planner, this allows real-time computation and safe, non-stop execution of multi-robot plans. The contributions of this work are:
  • a hierarchical framework for large-scale multi-robot real-time coordination that significantly reduces computation time compared to the baseline MAPF solver, while resulting in a moderately suboptimal solution; and
  • novel multi-commodity flow-based high-level planners, MCF/OD and one-shot MCF, that reduce congestion by regulating the influx of robots to each cell.
We demonstrate the algorithm in simulation with up to $142$ robots and in physical robot experiments with $32$ nano-quadrotors in cluttered environments, shown in Fig. 1.
II Related Work
Centralized approaches to multi-robot planning [5, 6] face substantial computational challenges due to the theoretical hardness of the problem [7], prohibiting real-time replanning and scaling to many robot systems. Decentralized approaches, on the other hand, can result in deadlocks, livelocks, congestion, collision, and reduced efficiency. RLSS [8] uses fast single-robot planners but relies on additional optimization to provide collision-free trajectories. In cluttered environments, a buffered Voronoi cell based algorithm [9] can result in deadlocks, and an algorithm using relative safe flight corridor [10] leads to collisions. In the present work, we aim to address these problems and facilitate real-time MAPF at the discrete planning phase.
Search-based MAPF solvers generate high-quality, collision-free discrete paths but the complexity for optimal solutions scales exponentially with the number of agents [7]. Bounded-suboptimal algorithms have been proposed to overcome this complexity, but their poor scalability still prevents their application to real-time coordination for large teams. To address this, partition-based MAPF [11, 12] divides the workspace into smaller regions, reducing the agent number within each cell. However, since the high-level planners are either single-agent based [11] or solving a multi-commodity flow problem constrained on single-robot shortest paths [12], they lead to congestion and hard MAPF instances. Instead, we propose a novel inter-cell routing algorithm to distribute robots while maintaining bounded-suboptimality. Furthermore, the cell-crossing method in [12] is unsuitable for aerial vehicles due to energy consumption while waiting in place for the cell-crossing channel to clear. Our cell-crossing protocol addresses this drawback.
Figure 2: The user inputs a map, start, and goal locations. Our approach generates a geometric partition, distributes robots among cells (hierarchical planner), and coordinates them within each cell in parallel.
III Problem Formulation
Consider a time-varying number of homogeneous non-point robots operating in workspace $\mathcal{W}$, which is partitioned into a union of non-overlapping convex polytopic cells. Robots must reach specified individual goal positions, which change over time, while avoiding collisions with robots and obstacles and obeying maximum cell influx limits $\boldsymbol{\theta}$ (influx refers to the number of robots entering a workspace cell). A motivating scenario is a multi-UAV package delivery system where the number of UAVs entering some types of airspace must be limited, and UAVs exit or enter the workspace to charge or redeploy.
The problem above requires solving a flow problem that obeys cell influx limits while handling new goal positions and a varying number of robots. While that is the general problem we aim to solve, the present work addresses a critical subproblem: the underlying planner that is repeatedly called to safely and efficiently route the robots through $\mathcal{W}$.
Consider $N$ (fixed) homogeneous non-point robots operating in a workspace $\mathcal{W}$ that is partitioned into a union of non-overlapping convex polytopic cells. $\mathcal{W}$ contains obstacles represented as unions of convex polytopes $\mathcal{O}_{1},\cdots,\mathcal{O}_{N_{obs}}$. Let $\mathcal{R}_{\mathcal{E}}(\mathbf{p})$ be the convex set of points representing a robot at position $\mathbf{p}$, i.e., a robot-environment collision model. The free space is represented as $\mathcal{F}=\mathcal{W}\backslash(\bigcup_{h}\mathcal{O}_{h})\ominus\mathcal{R}_{\mathcal{E}}(\mathbf{0})$, where $\ominus$ denotes Minkowski difference. For each robot $r^{i}$ (superscript $i$ represents robot index) with initial position $\mathbf{s}^{i}\in\mathbb{R}^{3}$, find paths for all robots to their goal positions such that there are no collisions (e.g., between robots or between robots and $\bigcup_{h}\mathcal{O}_{h}$) and the total number of robots that enter each cell is less than its user-defined influx $\theta_{m}$ (influx limits can vary by cell).
IV Preliminaries
IV-A Multi-agent pathfinding (MAPF)
Consider an undirected graph $\mathcal{G}\!\!=\!\!\left(V,E\right)$, and $N$ agents. Each agent has a start $v^{i}_{s}\!\!\in\!\!V$ and goal $v^{i}_{g}\!\!\in\!\!V$ vertex. At each time step $k$, an agent can either move to a neighbor vertex $\left(u^{i}_{k},u^{i}_{k+1}\right)\!\!\in\!\!E$ or stay at its current vertex $u^{i}_{k+1}\!\!=\!\!u^{i}_{k}$, where $u^{i}_{k}\!\!\in\!\!V$ is the $k$-th vertex in $i$-th agent’s path. To respect vertex conflict constraints, no two agents can occupy the same vertex simultaneously, i.e., $\forall k,i\!\!\neq\!\!j\!\!:\!\!u^{i}_{k}\neq u^{j}_{k}$. To respect edge conflict constraints, no two agents can traverse the same edge in the opposite direction concurrently, i.e., $\forall k,i\!\!\neq\!\!j\!\!:\!\!u^{i}_{k}\neq u^{j}_{k+1}\!\vee\!u^{i}_{k+1}\!\!\neq\!\!u^{j}_{k}$. The objective is to find conflict-free paths $\mathcal{P}^{i}\!\!=\!\!\left[u^{i}_{0},\cdots,u^{i}_{T-1}\right]$, where $u^{i}_{0}\!\!=\!\!v^{i}_{s}$ and $u^{i}_{T-1}\!\!=\!\!v^{i}_{g}$ for all agents, and minimize cost, e.g., the sum over the time steps required to reach the goals of all agents or the makespan $T$.
IV-B Conflict Annotation and MAPF with General Conflicts
Many works address MAPF for embodied agents [13, 14, 15]. We adopt multi-agent pathfinding with generalized conflicts (MAPF/C), due to its flexibility in incorporating different geometric shapes [5]. Conflict annotation identifies the extended conflict set of vertices and edges for each vertex and edge with respect to the inter-robot collision model. To account for the downwash effect between robots [16], we use a robot-robot collision model $\mathcal{R}_{\mathcal{R}}(\mathbf{p})$ that is distinct from $\mathcal{R}_{\mathcal{E}}(\mathbf{p})$. We follow the MAPF/C function definitions [5]:
$\operatorname{conVV}(v)=$$\left\{u\in V\mid\right.$
$\left.\mathcal{R}_{\mathcal{R}}(\operatorname{pos}(u))\cap\mathcal{R}_{\mathcal{R}}(\operatorname{pos}(v))\neq\emptyset\right\}$
$\operatorname{conEE}(e)=$$\left\{d\in E\mid\mathcal{R}_{\mathcal{R}}^{*}(d)\cap\mathcal{R}_{\mathcal{R}}^{*}(e)\neq\emptyset\right\}$
$\operatorname{conEV}(e)=$$\left\{u\in V\mid\right.\left.\mathcal{R}_{\mathcal{R}}(\operatorname{pos}(u))\cap\mathcal{R}_{\mathcal{R}}^{*}(e)\neq\emptyset\right\},$
where $\operatorname{pos}(u)\in\mathbb{R}^{3}$ returns the position for vertex $u$. $\mathcal{R}_{\mathcal{R}}^{*}(e)$ is the set of points swept by the robot when traversing edge $e$. This is called the swept collision model.
V Geometric Partitioning
Our framework has three components: geometric partitioning, high-level planner, and low-level planner, as depicted in Fig. 2. Geometric partitioning divides the workspace into disjoint convex cells, where the plan can be computed in parallel. A centralized high-level planner regulates the congestion for each cell while guaranteeing robots’ inter-cell routing quality. An anytime low-level planner plans collision-free paths for robots within each cell. Initial planning includes pre-computation (the geometric partitioning), and replanning only involves high- and low-level planning. Once the initial plan is established, our algorithm runs in real-time for replanning.
The geometric partitioning of a bounded workspace consists of three steps: 1) roadmap generation, 2) graph partitioning and spatial linear separation, and 3) local goal generation.
V-A Roadmap Generation
A roadmap is an undirected graph $\mathcal{G}\!\!=\!\!(V,E)$ embedded into a Euclidean space, where each vertex $v\!\!\in\!\!V$ corresponds to a position in $\mathcal{F}$ and each edge $(u,v)\!\!\in\!\!E$ denotes a path in $\mathcal{F}$ connecting $u$ and $v$. We additionally require the existence of vertices $v^{i}_{g}$ and $v^{i}_{s}$, corresponding to the goal and start positions, $\mathbf{g}^{i}$ and $\mathbf{s}^{i}$, respectfully. The roadmap should satisfy three properties: 1) connectivity-preserving, i.e., if a path between two points in $\mathcal{F}$ exists, there should be a path in the roadmap as well; 2) optimality-preserving, i.e., the shortest path between two points in $\mathcal{F}$ can be well approximated by a path in the roadmap; and 3) sparse, i.e., have a small number of vertices and edges. In our experiments, we use a 6-connected grid graph as roadmap, however, it can be generated by other methods, such as SPARS [17].
V-B Graph Partitioning and Spatial Linear Separation
Our method partitions the workspace into disjoint convex cells and within each cell solves a MAPF instance. Spatial decomposition has two benefits: 1) fewer robots for each smaller sub-graph, and 2) the decomposed MAPF instances can be solved in parallel. The initial partitioning can adopt any approach. We propose a geometric method that generates $Q$ convex polytopes. First, we use graph partitioning (KaHyPar [18]) to group the roadmap into $Q$ balanced sub-graphs $\mathcal{G}_{m}\!\!=\!\!\left({V_{m},E_{m}}\right)$, for $m\!\!=\!\!1,\ldots,Q$. Balanced means the vertex number difference between sub-graphs is bounded, leading to cells of similar volume and an even spread of robots.
Despite the sub-graphs now being balanced, we further enforce each cell, containing a sub-graph, to be a convex polytope. To do so, we use soft-margin support vector machines (SVM) [19] to compute the separating hyperplane $H_{ml}$ between sets of vertices of $\mathcal{G}_{m}$ and $\mathcal{G}_{l}$, and reassign misclassified vertices. The resulting set of separating hyperplanes form a cell. Convexity prevents robots from penetrating into neighboring regions while traversing within a cell, thus isolating cell planning.
(a)Refer to caption
(b)Refer to caption
(c)Refer to caption
(d)Refer to caption
Figure 3: (a) inter-robot collision configuration $\mathcal{C}_{col}$ and buffered hyperplanes. (b)-(d) vertex-vertex, edge-edge, and edge-vertex generalized conflicts across cells.
V-C Local Goal Generation
For navigating out of a cell, we generate candidate goal states on the faces between adjacent cells. For each cell $P_{m}$, we uniformly sample random local goals on the hyperplane $H_{ml}$ and add them as shared vertices to both $\mathcal{G}_{m}$ and $\mathcal{G}_{l}$. Despite being shared vertices, local goals generated by partition $P_{m}$ have in-edges from $P_{m}$ and out-edges to $P_{l}$ to avoid collision during cell transit. Thus, this part of the graph is directed. To enable parallel computation, there must be no communication between cells. Thus, the cell roadmap $\mathcal{G}_{m}$ is modified such that the planned paths are collision-free when crossing cells without information exchange between cells. Thus, the following properties should be satisfied:
P1: Robots avoid collision when stationary at vertices (local goals and non-local-goal vertices) of different cells (generalized vertex-vertex conflict across cells), i.e., $\forall i,j,m\!\!\neq\!\!l\!:\!v^{i}_{m}\!\!\notin\!\!\operatorname{conVV}(v^{j}_{l})$, where the superscript in the $v^{i}_{m}$ refers to the vertex index and the subscript refers to cell index.
P2: Robots avoid collision when traversing edges within different cells (generalized edge-edge conflict across cells), i.e., $\forall i,k,m\!\!\neq\!\!l:e^{i}_{m}\!\!:=\!\!(v^{i}_{m},v^{j}_{m})\notin\operatorname{conEE}(e^{k}_{l})$, here the superscript in $e^{i}_{m}$ refers to the edge index.
P3: Robots avoid collision when one robot is stationary at a vertex while the other robot is traversing an edge of a different cell (generalized edge-vertex conflict across cells), i.e., $\forall i,k,m\!\!\neq\!\!l:e^{i}_{m}:=(v^{i}_{m},v^{j}_{m}),v^{k}_{l}\notin\operatorname{conEV}(e^{i}_{m})$.
We depict violations of P1-P3 in Fig. 3b-3d. To prevent conflicts between local goals and non-local-goal vertices across cells, we buffer the separating hyperplane by the inter-robot collision configuration, $\mathcal{C}_{col}$ (c.f. Fig. 3a), which is computed as $\mathcal{C}_{col}(\mathbf{p})=\mathcal{R}_{\mathcal{R}}(\mathbf{p})\oplus\mathcal{R}_{\mathcal{R}}(\mathbf{0})$, where $\oplus$ is the Minkowski sum. The buffering is achieved by modifying the offset $\mathcal{H}_{a}^{{}^{\prime}}=\mathcal{H}_{a}+\mathrm{max}_{\mathbf{y}\in\mathcal{C}_{col}(\mathbf{0})}\mathcal{H}_{\mathbf{n}}\!\cdot\!\mathbf{y}$ of the hyperplane, where $\mathcal{H}_{a}$ and $\mathcal{H}_{\mathbf{n}}$ are the offset and the normal vector of the hyperplane. Given the buffered hyperplanes, all non-local-goal vertices within the buffered region are removed, guaranteeing collision avoidance between local goals and non-local-goal vertices. The local goals sampled on the hyperplane are confined within the blue region in Fig. 3a to avoid vertex-vertex conflicts between local goals of different separating hyperplanes. The buffering satisfies P1 between non-local-goal and local-goal vertices across cells and P1, P2, and P3 between non-local-goal vertices across cells. To satisfy P1 between local goals across cells, we uniformly randomly sample points on the hyperplane and reject those that violate P1. For P2 and P3 between non-local-goal and local goal vertices, we add the connection between the sampled local goal to its neighbor non-local-goal vertices within a radius on both sides of the hyperplane. The local goal is removed if any connected edge violates P2 or P3 (see Fig. 3c, 3d). Otherwise, the edges are added to the cell roadmaps. Figure 4 depicts an exemplar geometric partition.
(a)Refer to caption
(b)Refer to caption
Figure 4: Geometric partitioning with $Q\!\!=\!\!5$, before spatial linear separation (a), and after (b). Each cell is a convex polytope in the workspace.
VI Multi-commodity Flow with Optimal Detour
Our hierarchical approach relies on high-level planning to 1) regulate cell congestion and 2) preserve the bounded-suboptimality of inter-cell routing solutions. Thus, our high-level planner simplifies MAPF instances within cells and leads to real-time replanning. We abstract the partition as a directed graph $\mathcal{G}_{p}\!\!=\!\!\left(V_{p},E_{p}\right)$, where the vertices (nodes) represent cells and edges connect neighboring cells that share at least one face. Edges are weighted according to Euclidean distance between the cells’ centers of mass. The high-level planner finds a routing $\mathcal{U}^{i}\!\!=\!\!\left[P^{i}_{s},\cdots,P^{i}_{g}\right]$ for each robot $r^{i}$, where $P^{i}_{s}$ and $P^{i}_{g}$ are its start and goal cell, satisfying: 1) the influx of cell $m$ is under a user-defined value $\theta_{m}$, and 2) $cost(\mathcal{U}^{i})\!\!\leq\!\!w_{mcf}\cdot cost(\mathcal{U}^{i,*})$. Here, $w_{mcf}\!\!\geq\!\!1$ is a scalar representing the suboptimality bound for the routing solutions. $g^{i,*}$ is the optimal inter-cell routing of robot $r^{i}$. The influx of a node represents the number of robots entering the cell; we define influx formally in the following section.
VI-A High-level Planning Formulation
The SOTA partition-based MAPF solvers [11, 12] suffer from cell congestion, leading to hard MAPF instances in certain cells, causing computational bottlenecks. To address this, we formulate the inter-cell routing as a variant of the MCF problem and propose multi-commodity flow with optimal detour (MCF/OD), which optimally distributes robots among cells such that the number of robots entering any intermediate cell $m$ (not the start or goal cells of the robot teams) is under a user-defined value $\theta_{m}$, if solution exists.
Specifically, robots sharing the same start cell $P_{s}$ and goal cell $P_{g}$ are one commodity $c_{sg}$. The commodities set $C\!\!=\!\!\left\{c_{1},\cdots,c_{O}\right\}$ includes all commodities given the robot positions. Solving the MCF problem results in optimal flows $\{y^{*}_{sgml}\}$, that is the number of robots in commodity $c_{sg}\!\!\in\!\!C$ traversing along edge $e_{ml}\!\in\!E_{p}$. In our minimal influx MCF formulation, the optimal flow solutions lead to minimized intermediate cell influx and the most dispersed routing. The cell influx of $P_{l}$ is defined as the total number of robots entering this cell, that is, $\sum_{c_{sg},e_{ml}}y_{sgml}$. By constraining the flows $y_{sgml}$ to the shortest paths between the start and goal cells $SP_{sg}$, as in (1e), (1f), we guarantee solution optimality. We formulate the minimal influx MCF problem as the following integer linear program (ILP):
$\operatorname*{arg\,min}_{\left\{y_{sgml}\right\}}\quad\alpha\!$ $\cdot\!\sum_{c_{sg}}\mathcal{L}_{sg}+\beta\!\cdot\!\mathcal{L}_{in}$ (1a)
[…]
$\mathcal{L}_{sg}$ $\geq y_{sgml},~{}\forall e_{ml}\in E_{p},\forall c_{sg}\in C$ (1c)
$\mathcal{L}_{in}$ $\geq\mathbf{I}^{\top}_{v^{i}}\mathbf{y},~{}\forall v^{i}\in V_{p}$ (1d)
$y_{sgml}$ $\in\left[0,|c_{sg}|\right],~{}\forall e_{ml}\in SP_{sg},\forall c_{sg}\in C$ (1e)
$y_{sgml}$ $=0,~{}\forall e_{ml}\notin SP_{sg},\forall c_{sg}\in C\vspace{-0.5em}$ (1f)
where $\mathcal{L}_{sg}$ represents the maximum flow for commodity $c_{sg}\in C$, and $\mathcal{L}_{in}$ represents the maximum influx among all the cells. By minimizing both, the objective function penalizes the maximum congestion among all cells and disperses the flows related to one commodity. We weight $\mathcal{L}_{sg}$ and $\mathcal{L}_{in}$ with coefficients $\alpha$ and $\beta$, respectively, and set $\beta\gg\alpha$ to prioritize minimizing $\mathcal{L}_{in}$. In (1d), $\mathbf{I}_{v^{i}}=\left[I_{y_{1}},\cdots,I_{y_{F}}\right]$, where $y_{f}\in\left\{y_{sgml}\right\},I_{y_{f}}\in\left\{0,1\right\}$ is an indicator function that returns $1$ when $y_{f}$ has positive flow to an intermediate vertex $v^{i}\in V_{p}$. $\mathbf{y}$ is the vector of all flows.
Despite the minimized influx to the intermediate cells, the minimal influx MCF formulation leads to congestion in certain cells due to tight constraints on the shortest paths (robots’ shortest paths may intersect at certain cells). To detour the robots optimally, ensuring that the number of robots entering any cell $m$ remains below its influx limit $\theta_{m}$, we present a complete and optimal solver, MCF/OD, in Algo. 1. It maintains a conflict tree and resolves congestion iteratively. The function $congestionDetection(\mathcal{G}_{p},P.solution,\boldsymbol{\theta})$ computes the influx for each cell in $\mathcal{G}_{p}$, given the flow solution $P.solution$ and returns the set of cells with their influx larger than the corresponding limit. The function $getAllConflict(P)$ returns the set of all commodities passing congested cells.
1
Input: a multi-commodity flow instance.
2 Root.conflict_counts$[o]$ = $1,\text{for }o=1,\cdots,O$
3 Root.shortest_paths = solve shortest path for each commodity in the $\mathcal{G}_{p}$
4 Root.cost = sum of the largest cost among all paths in each commodity
5 Root.solution = solve MCF with root shortest path constraints
6 insert Root to OPEN
7 Visited[Root.conflict_counts] = True
8 while OPEN not empty do
9       $P\leftarrow$ node from OPEN with the lowest cost
10       $CS\leftarrow congestionDetection(\mathcal{G}_{p},P.solution,\boldsymbol{\theta})$
11       if CS is empty then
12             return P.solution
13      $C\leftarrow$ $getAllConflict\left(P,CS\right)$.
14       for commodity $c_{o}\in C$ do
15             $A\leftarrow$ new node.
16             $A$.conflict_counts $\leftarrow P$.conflict_counts + $\mathbf{I}_{o}$
17             if Visited[A.conflict_counts] then
18                   Continue
19            $A$.shortest_paths $\leftarrow P$.shortest_paths
[…]
21             Update $A$.shortest_paths[o] with $p_{k}$
22             if $p_{k}\leq w_{mcf}\cdot p_{min}$ then
23                   Update $A$.cost
24                   Update $A$.solution by solving MCF with updated shortest paths
[…]
26                         Insert $A$ to OPEN
27
28
29
Algorithm 1 MCF/OD
Theorem 1.
MCF/OD is complete on a locally finite graph.
Proof.
The cost of a conflict tree node equals the sum of the costs of the longest routing (without cycles) in all commodities. For each expansion, $k$-th shortest paths will be added to the commodity, which means the cost of the conflict tree is monotonically non-decreasing. For each pair of costs $X<Y$, the search will expand all nodes with cost $X$ before it expands the node with cost $Y$. As the graph is locally finite, there are a finite number of routing with the same cost for each commodity. Thus, expanding nodes with cost $X$ requires a finite number of iterations.
To include an arbitrary combination of $\hat{Z}$ unique edges of all commodities, the minimal cost of the conflict tree node is $Z$. $Z$ is finite, as the worst-case scenario is to include all the cell routing within the suboptimality bound. Since we are considering a graph with well-defined edge weights and a finite number of commodities, the worst cost is finite. For a finite cost $Z$, because the conflict tree node cost is monotonically non-decreasing and only a finite number of nodes with the same cost exists, we can find arbitrary combinations of $\hat{Z}$ unique edges in finite expansions. Thus, if a solution exists by including a combination of $\hat{Z}$ unique edges in the MCF, the algorithm can find it within finite expansions. If all the unique edges have been added to the MCF solver and the optimization cannot find the solution that satisfies the user-defined influx limit, the problem is identified as unsolvable. ∎
Theorem 2.
MCF/OD is optimal. If a solution is found, it will have the lowest possible cost, i.e., the sum of the costs of the longest routing in all commodities will be minimized if a solution is found.
Proof.
MCF/OD is a best-first search. In each expansion, the $k$-th shortest path for the selected commodity is inserted. Thus, the cost of a descendant node is monotonically non-decreasing. Therefore, if a solution is found, it is the optimal solution w.r.t. the cost. ∎
MCF/OD can find the optimal detouring solution. However, the complexity is high due to solving an ILP in each expansion. To tackle many commodities in a large partition, we propose another efficient detour algorithm, one-shot MCF, which solves MCF once. One-shot MCF augments the shortest paths in (1e), (1f) to include all the $w_{\mathrm{mcf}}$ bounded-suboptimal paths for each commodity. We employ the $k$-th shortest path routing algorithm to find all the candidate paths. The proposed One-shot MCF is complete as it includes all bounded-suboptimal paths for each commodity. While it does not optimize for the routing length for all commodities, it optimizes for minimum influx. Intuitively, MCF/OD adds bounded-suboptimal paths iteratively to relax the constraints and terminates once all cell influx limits are satisfied. On the other hand, One-shot MCF adds all bounded-suboptimal paths at once and optimizes for the minimum influx, so it could result in unnecessary detour.
In each high-level planning iteration, we run both MCF/OD and One-shot MCF in parallel. If MCF/OD times out, we use the solution generated from One-shot MCF. The high-level replanning happens every $\delta_{h}$ time interval.
VII Low-level Planner
Within each cell, the low-level planner, or the cell planner, computes collision-free paths that navigate robots to their local goals in an anytime fashion. The cell planner can be divided into three steps: 1) local goal assignment to determine the goal for each robot in a cell, 2) anytime MAPF/C to generate discrete paths, and 3) cell-crossing protocol for non-stop transiting between cells.
VII-A Local Goal Selection
At the beginning of cell planning, local goal selection aims to assign robots to the closest local goals while spreading out robots in an optimal manner by solving the following ILP:
$\operatorname*{arg\,min}_{\boldsymbol{\mathbf{A}}}\quad\sum_{ij}A_{ij}$ $\cdot D_{ij}+\alpha\sum_{j}u_{j}+\beta U$ (2a)
[…]
$U\geq u_{j}$ $\geq\sum_{i}A_{ij}-1,~{}\forall j,$ (2c)
where $A_{ij}\!\!\in\!\!\{0,1\}$ indicates if robot $r^{i}$ is assigned to local goal $lg^{j}$, and $D_{ij}$ is the Euclidean distance between $r^{i}$ and $lg^{j}$. Auxiliary variables $u_{j}$ in the objective function minimize the number of robots queueing at local goal $lg^{j}$, prioritizing filling less congested local goals first. Auxiliary variable $U$ in the objective function minimizes the maximum number of robots waiting in queue among all the local goals. This leads to evenly routing robots to different local goals to reduce congestion. A local goal is occupied if assigned with at least one robot. Thus, the number of robots waiting in queue for a local goal $lg^{j}$ is $\left(\sum_{i}A_{ij}-1\right)$.
VII-B Anytime MAPF/C
We adopt anytime MAPF, which iteratively improves solution quality until a solution is needed, to facilitate real-time replanning. We use the current SOTA anytime MAPF LNS [20], which iteratively replans for a subset of robots and adapts to an improved solution if found. We use ECBS [21] as the initial planner as it provides a bounded-suboptimal solution and prioritized planning with SIPP [22] to rapidly iterate plans. We extend both ECBS and SIPP using MAPF/C to account for generalized conflicts.
For priority planning with SIPP, we propose the following SIPP with generalized conflicts algorithm.
VII-B1 SIPP with generalized conflicts
SIPP compresses the time dimension into sparse safety intervals to significantly reduce the search space. The SIPP configuration augments the position with its safety intervals. In the resulting configuration space, $A^{*}$ finds the shortest path for a robot. Planned robots are considered moving obstacles and modify the safety interval of the traversed states. We propose SIPP with generalized conflicts (SIPP/C) with the following modifications and a different getSuccessors(s) algorithm, where the highlighted part differs from the original algorithm. In SIPP/C, the collision intervals, the complements of safety intervals, are added to vertices and edges based on the following vertex-vertex, edge-edge, and vertex-edge conflicts:
SIPP/C vertex-vertex conflict: for a robot at the vertex $u_{k}$ in a planned path at time $k$, we add the collision interval $[k,k]$ to vertices and edges that intersect with the robot-robot collision model at $pos(u_{k})$, i.e., $\mathcal{R}_{\mathcal{R}}(pos(v))\cap\mathcal{R}_{\mathcal{R}}(pos(u_{k}))$ and $\mathcal{R}^{*}_{\mathcal{R}}(e)\cap\mathcal{R}_{\mathcal{R}}(pos(u_{k}))$, $\forall v,e$. Note here, we use the swept model for $\mathcal{R}^{*}_{\mathcal{R}}(e)$ in SIPP.
SIPP/C edge-edge conflict: for a robot traversing an edge $e_{k}\!\!:=\!\!(u_{k},u_{k+1})$ at time $k$, we add the collision interval $[k,k]$ to vertices and edges that intersect with the robot’s swept collision model along edge $e_{k}$, i.e., $\mathcal{R}_{\mathcal{R}}(pos(v))\cap\mathcal{R}^{*}_{\mathcal{R}}(e_{k})$ and $\mathcal{R}^{*}_{\mathcal{R}}(e)\cap\mathcal{R}^{*}_{\mathcal{R}}(e_{k})$, $\forall v,e$.
SIPP/C edge-vertex conflict: for a robot to be stationary at the vertex $u_{k}\!\!=\!\!u_{k+1}$, we add the collision interval $[k,k]$ to vertices and edges that intersect with the robot’s swept collision model along edge $e_{k}\!\!:=\!\!(u_{k},u_{k+1})$, i.e., $\mathcal{R}_{\mathcal{R}}(pos(v))\cap\mathcal{R}_{\mathcal{R}}(pos(u_{k}))$ and $\mathcal{R}^{*}_{\mathcal{R}}(e)\cap\mathcal{R}_{\mathcal{R}}(pos(u_{k}))$, $\forall v,e$. Note that this degenerates to vertex-vertex conflict.
1 $successors=\emptyset$
2 for each action e in E(s) do
3       $cfg$ = configuration of $e$ applied to $s$
4       e_time = time to execute action $e$
5       start_t = time($s$) + e_time
6       end_t = interval($s$).end + e_time
7       for each safe interval i in cfg do
[…]
9                   continue
10            $t$ = earliest arrival time at $cfg$ during interval $i$ with no collisions
11             if interval($e$).end $&&\geq t\!-\!e\_time~{}\&\!\&$ interval($e$).start $\leq$ interval($s$).end &&~{}\&\!\&& & $i$.start $\leq$ interval($e$).start + e_time $\leq$ $i$.end  then
12                   t = max(interval($e$).start + m_time, t)
13
14            else
15                   continue
16             $s^{{}^{\prime}}$ = state of configuration $cfg$ with interval $i$ and time $t$
17             insert $s^{{}^{\prime}}$ into $successors$
18
return successors
Algorithm 2 getSuccessors(s) for SIPP/C
In Algo. 2, $E(s)$ is the action space at state $s$. Given the discrete path of robot $r^{i}$, we assign a time $t_{k}=k\Delta t$ to each discrete timestep and obtain the path $f^{i}$. $\Delta t$ is a user-defined value to satisfy the robot’s dynamic constraints. The low-level replanning happens every $\delta_{l}$ time interval.
VII-C Cell-crossing Protocol
A robot idles at a local goal if the path in the next cell is not yet computed. To allow non-stop execution, we propose a cell-crossing protocol that results in robots always having a planned path to execute, even when traversing between cells. We buffer the hyperplane $H_{ml}$ by a distance $d_{e}$ towards cell $P_{m}$. All robots within the buffer will compute paths for their next cell before leaving the current one. Buffering is achieved by changing the hyperplane offset to $\mathcal{H}^{{}^{\prime}}_{a}=\mathcal{H}_{a}-\mathcal{H}_{\mathbf{n}}\cdot d_{e}$. By enforcing the buffer distance $d_{e}\geq\delta_{l}\cdot V_{max}$, where $V_{max}$ is the maximum speed of the robot, the robot is guaranteed to have a plan computed at least once before leaving its current cell. A robot entering the buffer zone will then have a plan to exit its current cell and transition through its next cell. The robot concurrently computes plans for its current and next cell, then concatenates them to form a complete transition plan. The robot then fixes this plan to lock the local goal and expected arrival time to next cell. Thus, when computing the plan for the next cell, the robot’s expected start time and position will be pre-determined and independent of cell planning order. Fig. 5 depicts our cell-crossing protocol.
Figure 5: Robots (cones) fix their paths within the buffer zone and start replanning after crossing the hyperplane.
Instance Q Roadmap High-level Low-level
[…]
Demo32 8 1059 5890 MCF $0.06$ 0.06 18.0 ECBS(3.0)+PP $26.88$ $0.03\pm 0.01$ $0.13\pm 0.06$
Circle74 1 1110 6728 - - - 29.0 ECBS(2.0) $\mathbf{29.00}$ $12.31\pm 0.17$
Circle74 10 1515 8760 Greedy - - 42.4 ECBS(2.0)+PP $46.85$ $0.04\pm 0.03$ $0.40\pm 0.42$
Circle74 10 1515 8760 MCF $0.01$ 0.04 $\mathbf{23.9}$ ECBS(2.0)+PP $42.85$ $\mathbf{0.02\pm 0.00}$ $\mathbf{0.09\pm 0.02}$
Circle142 1 2100 13220 - - - 32.0 ECBS(4.5) $\mathbf{36.00}$ $16.32\pm 0.28$
Circle142 25 2954 16383 Greedy - - 73.6 ECBS(4.5)+PP $121.60$ $1.69\pm 0.82$ $37.58\pm 23.07$
Circle142 25 2954 16383 MCF $0.13$ 0.35 $\mathbf{27.1}$ ECBS(4.5)+PP $98.35$ $\mathbf{0.03\pm 0.01}$ $\mathbf{0.28\pm 0.14}$
TABLE I: Influence of different parameters on different instances. The statistics are averaged across 20 trials.
VIII Results and Discussion
[…]
Fig. 6 depicts typical solutions of the proposed algorithm. We summarize quantitative results in Table I, where MCF refers to MCF/OD and One-shot MCF running in parallel, as described in Sec. VI-A. Note that, as $\left|V\right|$ and $\left|E\right|$ suggest, the number of vertices and edges increases after partitioning as we add local goals and corresponding edges. With a small computational overhead $\bar{t}_{\mathrm{high}}$, the proposed high-level planner effectively reduces the congestion among cells compared to both greedy and partitionless baseline approaches, by inspecting $\bar{N}_{\mathrm{max}}$. Here $\bar{t}_{\mathrm{high}}$ is the average high-level planning time and $\bar{N}_{\mathrm{max}}$ is the averaged maximum number of robots in a cell throughout the whole execution. By increasing the number of cells $Q$, the algorithm significantly reduces MAPF computation time. Specifically, for the average low-level planning time $\bar{t}_{\mathrm{low}}$, in instance “Circle74”, the proposed algorithm runs $616$-times faster than the baseline method and $544$-times faster in instance “Circle142”. Both the average low-level replanning time $\bar{t}_{\mathrm{low}}$ and the averaged maximum low-level replanning time $\bar{t}^{\mathrm{max}}_{\mathrm{low}}$ are within the real-time regime for all instances. Since we use an anytime algorithm, we only record its initial planning time in the Table, which can be directly compared to the baseline. As we would expect, while performing in real-time, the algorithm yields suboptimal solutions compared to the baseline MAPF solution, according to the average makespan $\bar{T}$. Because the partitioning invalidates the global optimality of MAPF algorithms, and the high-level planner detours robots among the partition, lengthening planned paths.
(a)Refer to caption
(b)Refer to caption
(c)Refer to caption
Figure 6: Representative runs of our algorithm with (a) $74$ robots, (b) $142$ robots, and (c) $32$ robots in cluttered environments.
VIII-A Effectiveness of Partition in Low-level Planning
Fig. 7a shows a quantitative evaluation of the low-level planning time and its solution makespan when changing the number of cells in “Circle74”. As a robust statistic, we measure the median low-level planning time. As the result suggests, the logarithm of low-level planning time decreases significantly as we increase the number of cells. The result indicates that by partitioning the workspace and parallelizing computation, the low-level planning time decreases significantly. Furthermore, the makespan increases at the beginning then plateaus as the number of cells increases. This is expected since as the number of cells increases, first the high-level planner is more effective in detouring the robots to reduce congestion. At a certain point, no more detour is required since congestion is regulated. Additionally, the partition breaks the global optimality of the MAPF planner, leading to degeneracy in solution quality.
(a) Refer to caption
(b) Refer to caption
Figure 7: 95% Confidence intervals of (a) log-median of low-level planning time and Makespan of our algorithm (b) median of total low-level planning time. The statistics are averaged across $20$ trials.
Our algorithm runs MAPF in parallel for all cells; in Fig. 7b, we investigate the utilization of multi-threading. The orange dashed curve represents the theoretical lower bound of the total low-level planning time (assuming no overhead is introduced when allocating the computation to different threads). We observe the algorithm computation time generally follows the same curve, while the gap between the theoretical lower bound and experimental computation time increases as we increase the number of threads, indicating an increase in overhead for parallel computation in more cells.
VIII-B Effectiveness of High-level Planning
To demonstrate the effectiveness of our high-level planner, we compare the cell congestion of the proposed MCF-based algorithm to an egocentric greedy approach. The greedy planner outputs a single-robot-based shortest inter-cell routing without considering other robots’ routing, and may lead to congestion in certain partitions. Additionally, we compare the MCF-based approach with the baseline by imposing the partition into the workspace.
In Table. I, we report the average high-level computation time $\bar{t}_{\mathrm{high}}$ and the averaged maximum number of robots in a cell throughout the whole execution $\bar{N}_{\mathrm{max}}$. The complexity of a MAPF instance scales poorly with the number of robots. Thus, $\bar{N}_{\mathrm{max}}$ is a good indicator of the computational hardness of a MAPF instance. The computation time for greedy planning is instant while the MCF-based methods have additional overhead that increases with the number of cells, since they are centralized. For all instances in this paper, MCF-based methods provide real-time solutions.
The MCF-based methods reduce the congestion compared to greedy and baseline methods, according to $N_{\mathrm{max}}$. Although the advantage to the baseline is minor in the current experiment setup, the MCF-based approach can further reduce the congestion by relaxing bounded-suboptimality constraints, i.e., increase $w_{mcf}$. For a fair comparison of low-level planning time, both greedy and MCF-based utilize the same number of threads in computation. Compared to the greedy approach, the proposed MCF-based high-level planning yields more efficient low-level planning, which brings computation time to real time while maintaining the solution quality, i.e., makespan.
VIII-C Scalability of the Proposed Algorithm
To investigate the scalability of the proposed algorithm, we run simulations on increasing numbers of robots (c.f., Table. I). Note that the proposed algorithm achieves real-time performance in all instances as we scale up the number of robots, by looking at the averaged maximum high-level and low-level replanning time $\bar{t}^{\mathrm{max}}_{\mathrm{high}}$ and $\bar{t}^{\mathrm{max}}_{\mathrm{low}}$. Due to the hierarchical approach, the proposed algorithm can further scale to larger teams with more cells and CPU threads. The computation bottleneck is the centralized high-level planning in larger scale problems, which we aim to address in future work.
VIII-D Physical Robots
Fig. 1 shows a representative experiment with $32$ physical Crazyflies in a cluttered environment (video link: https://youtu.be/ftdWVpLkErs). We use a Vicon motion capture system to localize and CrazySwarm [23] to control the robots. In the experiment, initial positions are uniformly spaced on an ellipse within the workspace, and the goal positions are the antipodal points on the ellipse. Three column obstacles are placed within the ellipse. The experiment demonstrates that the proposed algorithm distributes robots effectively through the workspace and achieves real-time replanning.
IX Conclusion and Future Work
We have introduced a hierarchical path planning algorithm for large-scale coordination tasks. Our algorithm significantly reduces the computation time and suits on-demand applications such as drone delivery despite yielding a suboptimal solution. The framework achieves real-time operation by dividing the workspace into disjoint convex cells; within each, an anytime MAPF planner computes collision-free paths in parallel. Our high-level planner regulates the congestion of regions while guaranteeing the routing quality. Additionally, our algorithm considers the robot’s geometric shape constraints in continuous space, and we run experiments with the collision model of a quadrotor with downwash. We also devise a cell-crossing protocol, which guarantees the robot always has a plan, even when transiting between cells, and allows replanning in a continuous time domain.
The proposed algorithm is designed for lifelong replanning. In our experiments, goals are chosen from a pre-determined set, however, it can be extended to a general lifelong replanning as we add and delete goals online. When new goals are requested, the algorithm can identify the new goal’s cell, and run conflict annotation steps in parallel for each new goal.
While the MCF-based high-level planner operates in real-time in our experiments with up to $142$ robots, the limits of its real-time operation are not well defined and depend on the number and density of robots and cells in the space. Future work will explore distributed MCF [24, 25] to increase scalability of the system with real-time, distributed operation regardless of the number of robots and cells. Furthermore, we aim to solve real time large scale motion planning, which respects the robot’s kinodynamic constraints and plans in continuous space and time.
X Acknowledgements
The authors would like to thank Baskın Şenbaşlar, Eric Ewing, Yuan Yuan, Calvin Luo, Yutong Wang, Arjun Prakash for their help in revising the manuscript.